# Self-Speculative Masked Diffusions

### v1

@@ Appendix C Proofs
$p(\bm{x},A)=\text{min}(\overset{\leftrightarrow}{p}(\bm{x}),\overset{\rightarrow}{p}(\bm{x})),$ (26)
@@
$p(\bm{x},R)=\text{max}(0,\overset{\rightarrow}{p}(\bm{x})-\overset{\leftrightarrow}{p}(\bm{x}))$ (27)
@@
We introduce labels for all random variables involved in this procedure. We have the draft token $\hat{\bm{x}}$, a uniform random variable for making the accept/reject decision, $U\sim\mathcal{U}(0,1)$, the resampled token $\tilde{\bm{x}}$, the accept/reject decision variable $Y\in\{A,R\}$ and finally the output token $\bm{x}$.
We can write a joint distribution over all variables involved as
$p(\hat{\bm{x}},U,\tilde{\bm{x}},Y,\bm{x})=p(\hat{\bm{x}})p(U)p(\tilde{\bm{x}})p(Y|U,\hat{\bm{x}})p(\bm{x}|\hat{\bm{x}},\tilde{\bm{x}},Y)$ (28)
@@
$p(\bm{x}|\hat{\bm{x}},\tilde{\bm{x}},Y)$ $=\updelta\{\bm{x}=\hat{\bm{x}}\}\updelta\{Y=A\}+\updelta\{\bm{x}=\tilde{\bm{x}}\}\updelta\{Y=R\}$ (30)
where $\updelta$ is the Kronecker delta.
We first find $p(\bm{x},A)$ by summing/integrating out all other variables and setting $Y=A$,
$p(\bm{x},A)$ $=\sum_{\hat{\bm{x}}}\int_{U=0}^{U=1}\sum_{\tilde{\bm{x}}}p(\hat{\bm{x}},U,\tilde{\bm{x}},Y=A,\bm{x})\mathrm{d}U$ (31)
$=\sum_{\hat{\bm{x}}}\int_{U=0}^{U=1}\sum_{\tilde{\bm{x}}}\overset{\leftrightarrow}{p}(\hat{\bm{x}})p_{\text{res}}(\tilde{\bm{x}})\updelta\left\{U<\text{min}\left(1,\frac{\overset{\rightarrow}{p}(\hat{\bm{x}})}{\overset{\leftrightarrow}{p}(\hat{\bm{x}})}\right)\right\}\updelta\{\bm{x}=\hat{\bm{x}}\}\mathrm{d}U$ (32)
$=\sum_{\hat{\bm{x}}}\overset{\leftrightarrow}{p}(\hat{\bm{x}})\updelta\{\bm{x}=\hat{\bm{x}}\}\int_{U=0}^{U=1}\updelta\left\{U<\text{min}\left(1,\frac{\overset{\rightarrow}{p}(\hat{\bm{x}})}{\overset{\leftrightarrow}{p}(\hat{\bm{x}})}\right)\right\}\mathrm{d}U\underbrace{\sum_{\tilde{\bm{x}}}p_{\text{res}}(\tilde{\bm{x}})}_{=1}$ (33)
[…]
$=\overset{\leftrightarrow}{p}(\bm{x})\text{min}\left(1,\frac{\overset{\rightarrow}{p}(\bm{x})}{\overset{\leftrightarrow}{p}(\bm{x})}\right)$ (35)
$=\text{min}\left(\overset{\leftrightarrow}{p}(\bm{x}),\overset{\rightarrow}{p}(\bm{x})\right).$ (36)
We next find $p(\bm{x},R)$ again by summing/integrating out other variables and setting $Y=R$,
$p(\bm{x},R)$ $=\sum_{\hat{\bm{x}}}\int_{U=0}^{U=1}\sum_{\tilde{\bm{x}}}p(\hat{\bm{x}},U,\tilde{\bm{x}},Y=R,\bm{x})\mathrm{d}U$ (37)
$=\sum_{\hat{\bm{x}}}\int_{U=0}^{U=1}\sum_{\tilde{\bm{x}}}\overset{\leftrightarrow}{p}(\hat{\bm{x}})p_{\text{res}}(\tilde{\bm{x}})\delta\left\{U>\text{min}\left(1,\frac{\overset{\rightarrow}{p}(\hat{\bm{x}})}{\overset{\leftrightarrow}{p}(\hat{\bm{x}})}\right)\right\}\delta\{\bm{x}=\tilde{\bm{x}}\}\mathrm{d}U$ (38)
$=\sum_{\hat{\bm{x}}}\int_{U=0}^{U=1}\overset{\leftrightarrow}{p}(\hat{\bm{x}})p_{\text{res}}(\bm{x})\delta\left\{U>\text{min}\left(1,\frac{\overset{\rightarrow}{p}(\hat{\bm{x}})}{\overset{\leftrightarrow}{p}(\hat{\bm{x}})}\right)\right\}\mathrm{d}U$ (39)
[…]
$)\displaystyle=p_{\text{res}}(\bm{x})\sum_{\hat{\bm{x}}}\text{max}(0,\overset{\leftrightarrow}{p}(\hat{\bm{x}})-\overset{\rightarrow}{p}(\hat{\bm{x}}$)) (42)
$)\displaystyle=p_{\text{res}}(\bm{x})\sum_{\hat{\bm{x}}}\text{max}(0,\overset{\rightarrow}{p}(\hat{\bm{x})}-\overset{\leftrightarrow}{p}(\hat{\bm{x}}$)) (43)
$)\displaystyle=\frac{\text{max}(0,\overset{\rightarrow}{p}(\bm{x})-\overset{\leftrightarrow}{p}(\bm{x}))}{\sum_{z}\text{max}(0,\overset{\rightarrow}{p}(z)-\overset{\leftrightarrow}{p}(z))}\sum_{\hat{\bm{x}}}\text{max}(0,\overset{\rightarrow}{p}(\hat{\bm{x}})-\overset{\leftrightarrow}{p}(\hat{\bm{x}}$)) (44)
$=\text{max}(0,\overset{\rightarrow}{p}(\bm{x})-\overset{\leftrightarrow}{p}(\bm{x}))$ (45)
where in Equation 43 we have used the fact that $)\sum_{\hat{\bm{x}}}\text{max}(0,\overset{\leftrightarrow}{p}(\hat{\bm{x}})-\overset{\rightarrow}{p}(\hat{\bm{x}}))=\sum_{\hat{\bm{x}}}\text{max}(0,\overset{\rightarrow}{p}(\hat{\bm{x}})-\overset{\leftrightarrow}{p}(\hat{\bm{x}}$)). This can be seen through the following manipulations,
[…]
This completes the proof of Lemma C.1. ∎
@@ C.2 Number of Rejections
$=$ $\left\{\prod_{k=d}^{D}p_{\theta,\phi}(\bm{x}^{k},A^{k}|\bm{x}^{1:k-1},A^{d:k-1},R^{d-1},N^{d-1})\right\}$ (112)
$\times p_{\theta,\phi}(N^{D}|\bm{x}^{1:D},A^{d:D},R^{d-1},N^{d-1}).$ (113)
@@ Appendix D Choices of Window Function
In our experiments, however, we found that a cosine shaped window works the best, similar to the finding for the schedule in MDMs (Shi et al.,, 2024).
@@
$W(i)$ $=D\left(\cos\left(\frac{\pi}{2}\left(1-\tau\right)\right)-\cos\left(\frac{\pi}{2}\left(1-\tau+\Delta\tau\right)\right)\right).$ (147)
Appendix E Experimental Details
@@
E.1 Text8
For both the mask diffusion and speculative model, we use a 12 layer transformer with 12 heads and 768 hidden dimension following the architecture used by Shi et al., (2024). The model was trained for 1M steps with a batch size of 256. We use a learning rate schedule with $2000$ linear warm-up steps with a maximum learning rate value of $3\times 10^{-4}$. The learning rate is then decayed using a cosine schedule over the 1M training steps. We use a dropout rate of 0.05 within the transformer and apply weight decay with parameter $0.03$. For the mask diffusion baseline, we use a cosine noise schedule as used by Shi et al., (2024).
@@ E.1 Text8
When sampling the mask diffusion baseline, in order to avoid the truncation issue described in Zheng et al., (2025), we first sample $\bm{x}_{0}$ from the model’s denoising distribution and then randomly select some number of currently masked positions to switch value to the corresponding $\bm{x}_{0}$ value. The probability of selecting a masked position to be revealed is given by the noise schedule. This avoids computing the probability as a combination of both the reveal probability and the token value probability which results in the truncation problem.
@@
E.2 OpenWebText
For our mask diffusion baseline, we use the 12 block transformer from Shi et al., (2024). As our self-speculative model, we switch the last block to causal resulting in an 11 non-causal block 1 causal block architecture. Our training setup follows that of Shi et al., (2024), training for 1M iterations with a batch size of 256 on 64 TPUv3 chips. The training losses are given in Figure 6. Each element in the batch is a sequence of tokens of length 1024. We set the dropout rate to 0 and use a weight decay value of $0.03$. For our noising schedule during training, we use the cosine noise schedule following Shi et al., (2024). The learning rate has a 2000 step linear warmup to a maximum value of $3\times 10^{-4}$ which is then decayed using the cosine schedule.
@@ E.2 OpenWebText
For the SDTT baseline (Deschenaux and Gulcehre,, 2025), we run the author provided code available at https://github.com/jdeschena/sdtt. The authors find their KLD objective to perform the best with the maximum speed-up observed on the final round of self-distillation (7 rounds in their work). We therefore compute the baseline with the provided KLD round 7 model. We run the generated samples through our GPT2 NLL computation pipeline to ensure consistency in our comparison.
As for the text8 experiment, we again provide example outputs from our speculative model and the masked diffusion baseline for both low and high NFE in Section E.2.1.
@@ E.2.1 Samples
E.3 Uniref50
We build our experiments on top of the codebase provided by Wang et al., (2025), https://github.com/bytedance/dplm. We take the airkingbd/dplm_150m as our base masked diffusion model which is based on the 150M parameter ESM2 network (Lin et al.,, 2023). The standard network contains 30 non-causal blocks and we add an additional block with the same architecture as the previous 30 however switching to a causal attention mask. We interface the non-causal blocks and causal block using the wiring diagram given in Figure 1. We fine-tune the additional causal block for $500\times 10^{3}$ iterations, the training losses are shown in Figure 7. As the non-causal blocks are frozen, their average losses remain constant during fine-tuning, the noise being caused by different datapoints being sampled in each batch along with different masking ratios. We see that during fine-tuning the causal block learns to use the information provided in the extra tokens revealed to it in order to achieve a lower loss. Thus this provides a target distribution that better models the data than the non-causal draft distribution.
The base ESM2 network utilizes rotary position embeddings or RoPE (Su et al.,, 2024) to encode the sequence positions rather than position encodings that are concatenated to the input as shown in Figure 1. RoPE rotates the query and key values in self-attention by angles given by that track’s position in the sequence. We would like to use the double embedding system from $\sigma$-GPT with RoPE, where both the current position and the next position in the ordering is encoded. To achieve this we split the RoPE channels in half, with the first half encoding the current position and the second half encoding the next position in the sequence.

### v2

@@ Appendix C Proofs
$p(\bm{x},A)=\min(\overset{\leftrightarrow}{p}(\bm{x}),\overset{\rightarrow}{p}(\bm{x})),$ (26)
@@
$p(\bm{x},R)=\max(0,\overset{\rightarrow}{p}(\bm{x})-\overset{\leftrightarrow}{p}(\bm{x})).$ (27)
@@
$p(\bm{x},A)$ $=\overset{\leftrightarrow}{p}(\bm{x})\min\left(1,\frac{\overset{\rightarrow}{p}(\bm{x})}{\overset{\leftrightarrow}{p}(\bm{x})}\right)=\min(\overset{\leftrightarrow}{p}(\bm{x}),\overset{\rightarrow}{p}(\bm{x})).$ (28)
For a rejected and resampled token $\bm{x}$, it must be drawn from $p_{\text{res}}(\bm{x})$ after some draft token is rejected. The total probability of rejection, $P(R)$, is the expected rejection rate over all possible draft tokens:
@@
$)\displaystyle=\sum_{\hat{\bm{x}}}\max(0,\overset{\leftrightarrow}{p}(\hat{\bm{x}})-\overset{\rightarrow}{p}(\hat{\bm{x}}$)) (31)
$=\sum_{\hat{\bm{x}}}\max(0,\overset{\rightarrow}{p}(\hat{\bm{x}})-\overset{\leftrightarrow}{p}(\hat{\bm{x}}))\quad\text{as~}\sum_{\hat{\bm{x}}}(\overset{\rightarrow}{p}(\hat{\bm{x}})-\overset{\leftrightarrow}{p}(\hat{\bm{x}}))=0.$ (32)
Finally, given $p(\bm{x},R)=P(R)\cdot p_{\text{res}}(\bm{x})$, we obtain
$p(\bm{x},R)=P(R)\cdot\frac{\max(0,\overset{\rightarrow}{p}(\bm{x})-\overset{\leftrightarrow}{p}(\bm{x}))}{P(R)}=\max(0,\overset{\rightarrow}{p}(\bm{x})-\overset{\leftrightarrow}{p}(\bm{x})).$ (33)
This completes the proof. ∎
@@ C.2 Number of Rejections
$=$ $\left\{\prod_{k=d}^{D}p_{\theta,\phi}(\bm{x}^{k},A^{k}|\bm{x}^{1:k-1},A^{d:k-1},R^{d-1},N^{d-1})\right\}p_{\theta,\phi}(N^{D}|\bm{x}^{1:D},A^{d:D},R^{d-1},N^{d-1}).$ (95)
@@ Appendix D Choices of Window Function
In our experiments, however, we found that a cosine shaped window works the best, similar to the finding for the schedule in MDMs (Shi et al., 2024).
@@
$W(i)$ $=\left\lfloor D\left(\cos\left(\frac{\pi}{2}\left(1-\tau\right)\right)-\cos\left(\frac{\pi}{2}\left(1-\tau+\Delta\tau\right)\right)\right)\right\rfloor,$ (129)
where $\lfloor\cdot\rfloor$ is the floor function.
Appendix E FLOP Analysis
Our analysis follows that of Hoffmann et al. (2022) Appendix F. The vanilla result holds for both autoregressive models and MDMs since these two differ only in the attention mask which does not affect the total FLOPs.
We make the following definitions: let $C$ be the base hidden dimension, $F$ be the feed-forward hidden dimension, $H$ be the number of heads, $K$ be the number of dimensions in the key, $V$ be the vocabulary size and $S$ be the sequence length. Hoffmann et al. (2022) then provide the following relations for the number of FLOPs in each part of a transformer in terms of $C,F,H,K,V,S$.
Embedding $=2SVC$
Single layer attention:
$\text{QKV projection}=6SCKH$
$\text{K@Q}=2S^{2}KH$
$\text{Softmax}=3HS^{2}$
$\text{Softmax @ query reduction}=2S^{2}KH$
$\text{Linear}=2SKHC$
Dense block $=4SCF$
Final logits $=2SCV$
The total FLOPs for a forward pass of a vanilla transformer network is then Total vanilla FLOPs = embedding + num layers * (attention + dense) + logits.
In our self-speculative architecture, the FLOPs calculation within the blocks are unchanged however we additionally include a projection step before the causal block to include the current and next state information. We re-use the input token embeddings from the standard input to the transformer and in our experiments we used rotary positional encodings to encode the position information, splitting the channel dimension in half between the current position and next position. Therefore the only extra FLOPs incurred at this stage are those projecting the concatenated current position hidden state, next position hidden state and token embedding.
A projection operation is of the form $h=Wx$ where $x\in\mathbb{R}^{c_{in}}$, $W\in\mathbb{R}^{c\times c_{in}}$ and $h\in\mathbb{R}^{c}$. To compute the matrix multiplication there are $c_{in}$ multiplications and $c_{in}$ additions for each element in $h$ meaning a total FLOP count of $2c_{in}c$. In our case, we concatenate two hidden states together and a token embedding and project back down to $C$ giving a FLOP count of $2\times 3C\times C=6C^{2}$ per token. In addition, we add on the non-causal hidden states to the output causal hidden states before projecting them to the causal distribution logits which is $C$ add operations, resulting in an extra $C$ FLOPs per token. For the architecture settings used in our experiments, we can now compare the FLOPs for the vanilla architecture versus the extra overhead FLOPs incurred by our method. Our transformer on OpenWebText used the following values: $C=768$, $V=50,257$, $K=64$, $H=12$, $F=3072$, $S=1024$, $\text{num layers}=12$. The vanilla transformer FLOPs are then
Embedding $=7.9e10$
Single layer attention:
$\text{QKV projection}=3.6e9$
$\text{K@Q}=1.6e9$
$\text{Softmax}=3.7e7$
$\text{Softmax @ query reduction}=1.6e9$
$\text{Linear}=1.2e9$
$\text{Total}=8e9$
Dense block $=9.7e9$
Final logits $=7.9e10$
Total vanilla FLOPs $=7.9e10+12\times(8e9+9.7e9)+7.9e10=3.7e11$
The additional overhead FLOPs in our method are $S\times(6C^{2}+C)=3.6e9$. Therefore the overhead FLOPs are only 0.98% of the value of the total FLOPs for a forward pass of the entire network and so are insignificant when compared to the 2x reduction in total number of forward passes required to generate a datapoint.
Appendix F Hyperparameter Influence on Performance
In this section we discuss the influence of the hyperparameters of our overall self-speculative approach on sample quality and inference time.
In terms of the number of non-causal vs causal layers, Table 1 finds that if the number of causal blocks is increased to 2 and non-causal blocks decreased to 10 then we find a worse trade-off between sample quality and NFE likely due to stronger target not making up for the weaker draft model. We expect this trend to continue to higher numbers of causal blocks.
For the number of verification steps per draft and the $\Delta\tau$ parameter of the cosine window, we found in Figure 3 that increasing $\Delta\tau$ results in more tokens being accepted per draft step. However, this increased limit on the number of revealed tokens can result in occasions where too many tokens are revealed from a weak target early in generation causing worse generation quality. We can see this by focusing on the spelling accuracy - NFE tradeoff in the regime where $\Delta\tau$ is varied and the number of verification steps per draft is held constant at 1.
$\Delta\tau$ Spelling Accuracy NFE
0.01 0.91 80
0.02 0.90 44
0.04 0.88 28
0.083 0.87 21
Table 2: Table highlighting the specific influence of $\Delta\tau$ on spelling accuracy on the text8 experiment in Section 5.1.
We observe that initially as $\Delta\tau$ is increased away from 0.01, there is a large reduction in NFE with only a marginal decrease in spelling accuracy however, the spelling accuracy hit becomes worse as $\Delta\tau$ is increased further and more tokens are allowed to be revealed per step.
We also find in Figure 3 that increasing the number of verification steps per draft has a similar effect to $\Delta\tau$ in that it allows more tokens to be revealed for each pass of the non-causal blocks which can decrease sample quality due to more tokens being revealed at the start of sampling when the model has less information on the final datapoint resulting in a weaker predictive distribution.
The ultimate hyperparameter selection that a practitioner should use will depend on their use case and their requirements on sample quality and latency. Should latency be a major factor, then $\Delta\tau$ and the number of verification steps per draft should be increased up until the point that sample quality becomes unsatisfactorily low.
Appendix G Experimental Details
@@
G.1 Text8
For both the mask diffusion and speculative model, we use a 12 layer transformer with 12 heads and 768 hidden dimension following the architecture used by Shi et al. (2024). The model was trained for 1M steps with a batch size of 256. We use a learning rate schedule with $2000$ linear warm-up steps with a maximum learning rate value of $3\times 10^{-4}$. The learning rate is then decayed using a cosine schedule over the 1M training steps. We use a dropout rate of 0.05 within the transformer and apply weight decay with parameter $0.03$. For the mask diffusion baseline, we use a cosine noise schedule as used by Shi et al. (2024).
@@ G.1 Text8
When sampling the mask diffusion baseline, in order to avoid the truncation issue described in Zheng et al. (2025), we first sample $\bm{x}_{0}$ from the model’s denoising distribution and then randomly select some number of currently masked positions to switch value to the corresponding $\bm{x}_{0}$ value. The probability of selecting a masked position to be revealed is given by the noise schedule. This avoids computing the probability as a combination of both the reveal probability and the token value probability which results in the truncation problem.
@@
G.2 OpenWebText
For our mask diffusion baseline, we use the 12 block transformer from Shi et al. (2024). As our self-speculative model, we switch the last block to causal resulting in an 11 non-causal block 1 causal block architecture. Our training setup follows that of Shi et al. (2024), training for 1M iterations with a batch size of 256 on 64 TPUv3 chips. The training losses are given in Figure 6. Each element in the batch is a sequence of tokens of length 1024. We set the dropout rate to 0 and use a weight decay value of $0.03$. For our noising schedule during training, we use the cosine noise schedule following Shi et al. (2024). The learning rate has a 2000 step linear warmup to a maximum value of $3\times 10^{-4}$ which is then decayed using the cosine schedule.
@@ G.2 OpenWebText
For the SDTT baseline (Deschenaux and Gulcehre, 2025), we run the author provided code available at https://github.com/jdeschena/sdtt. The authors find their KLD objective to perform the best with the maximum speed-up observed on the final round of self-distillation (7 rounds in their work). We therefore compute the baseline with the provided KLD round 7 model. We run the generated samples through our GPT2 NLL computation pipeline to ensure consistency in our comparison.
As for the text8 experiment, we again provide example outputs from our speculative model and the masked diffusion baseline for both low and high NFE in Section G.2.1.
@@ G.2.1 Samples
G.3 Uniref50
We build our experiments on top of the codebase provided by Wang et al. (2025), https://github.com/bytedance/dplm. We take the airkingbd/dplm_150m as our base masked diffusion model which is based on the 150M parameter ESM2 network (Lin et al., 2023). The standard network contains 30 non-causal blocks and we add an additional block with the same architecture as the previous 30 however switching to a causal attention mask. We interface the non-causal blocks and causal block using the wiring diagram given in Figure 1. We fine-tune the additional causal block for $500\times 10^{3}$ iterations, the training losses are shown in Figure 7. As the non-causal blocks are frozen, their average losses remain constant during fine-tuning, the noise being caused by different datapoints being sampled in each batch along with different masking ratios. We see that during fine-tuning the causal block learns to use the information provided in the extra tokens revealed to it in order to achieve a lower loss. Thus this provides a target distribution that better models the data than the non-causal draft distribution.
The base ESM2 network utilizes rotary position embeddings or RoPE (Su et al., 2024) to encode the sequence positions rather than position encodings that are concatenated to the input as shown in Figure 1. RoPE rotates the query and key values in self-attention by angles given by that track’s position in the sequence. We would like to use the double embedding system from $\sigma$-GPT with RoPE, where both the current position and the next position in the ordering is encoded. To achieve this we split the RoPE channels in half, with the first half encoding the current position and the second half encoding the next position in the sequence.